# LLM Knows Geometry Better than Algebra: Numerical Understanding of LLM-Based Agents in A Trading Arena

### v1

@@ Abstract
Recent advancements in large language models (LLMs) have significantly improved performance in natural language processing tasks. However, their ability to generalize to dynamic, unseen tasks, particularly in numerical reasoning, remains a challenge. Existing benchmarks mainly evaluate LLMs on problems with predefined optimal solutions, which may not align with real-world scenarios where clear answers are absent. To bridge this gap, we design the Agent Trading Arena, a virtual numerical game simulating complex economic systems through zero-sum games, where agents invest in stock portfolios. Our experiments reveal that LLMs, including GPT-4o, struggle with algebraic reasoning when dealing with plain-text stock data, often focusing on local details rather than global trends. In contrast, LLMs perform significantly better with geometric reasoning when presented with visual data, such as scatter plots or K-line charts, suggesting that visual representations enhance numerical reasoning. This capability is further improved by incorporating the reflection module, which aids in the analysis and interpretation of complex data. We validate our findings on NASDAQ Stock dataset, where LLMs demonstrate stronger reasoning with visual data compared to text. Our code and data are publicly available at https://github.com/wekjsdvnm/Agent-Trading-Arena.git.
LLM Knows Geometry Better than Algebra:
Numerical Understanding of LLM-Based Agents in A Trading Arena
Tianmi Ma1, Jiawei Du3,4,, Wenxin Huang2, ✉, Wenjie Wang5, Liang Xie6, Xian Zhong1, ✉, and Joey Tianyi Zhou3,4 1 Hubei Key Laboratory of Transportation Internet of Things, Wuhan University of Technology 2 Hubei Key Laboratory of Big Data Intelligent Analysis and Application, Hubei University 3 Centre for Frontier AI Research, Agency for Science, Technology and Research, Singapore 4 Institute of High Performance Computing, Agency for Science, Technology and Research, Singapore 5 School of Computing, National University of Singapore  6 School of Science, Wuhan University of Technology
footnotetext: Email:wenxinhuang_wh@163.com, zhongx@whut.edu.cn. denotes the equal contribution. ✉ represents the corresponding author.
@@ 1 Introduction
Recent advancements in large language models (LLMs) have demonstrated exceptional proficiency across various domains, achieving state-of-the-art performance in natural language processing (NLP) tasks such as translation (Koshkin et al., 2024), summarization (Yang et al., 2023), and reasoning (Kalyanpur et al., 2024; Qiao et al., 2023). While LLMs excel in language-based tasks, further progress in numerical and geometric reasoning is essential to tackling complex, interdisciplinary challenges, particularly in fields like finance and scientific research. Benchmarks such as GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) have been developed to assess and improve the mathematical problem-solving abilities of LLMs through structured datasets and standardized evaluation protocols. These benchmarks not only advance LLM development but also represent a crucial step in enhancing the ability of artificial intelligence systems to solve mathematical reasoning problems.
Figure 1: Illustration and Performance Comparison of Textual and Visual (K-Line Chart) Numerical Inputs. Top: The left figure presents numerical data in K-line chart format, while the right figure shows textual numerical inputs. Bottom: LLMs with visual inputs (orange) significantly outperform those with textual inputs (blue) in 40-day yield trends, demonstrating the advantage of visual data in financial decision-making.
In response, recent updates to LLMs have focused on improving their mathematical abilities, with models (Ahn et al., 2024; Romera-Paredes et al., 2024) achieving new records on math-focused benchmarks. However, many of these benchmarks (Taylor et al., 2022; Naveed et al., 2023) primarily evaluate performance on problems familiar to the models, closely resembling those encountered during training and often following standard formats, solution strategies, or recurring mathematical patterns. This raises concerns about whether LLMs genuinely possess reasoning abilities for novel numerical data or whether their success is largely due to memorization and pattern recognition. These limitations underscore the need for alternative evaluation paradigms that assess LLMs’ ability to generalize mathematical principles to unseen scenarios.
To address these limitations, we introduce the Agent Trading Arena, a virtual numerical game generating diverse numerical data through agent interactions in a zero-sum stock market environment. In this system, LLM-based agents make trading decisions based on historical stock prices, dynamically determined by agents’ bidding activities. Unlike static benchmarks (Naveed et al., 2023) with predefined strategies, our simulation forces agents to adapt to evolving market conditions. Any temporarily optimal strategy is quickly countered, fostering an environment that rewards adaptability. By mimicking competitive trading scenarios where outcomes depend on real-time agent decisions, we create a controlled yet dynamic environment that tests LLMs’ ability to infer hidden patterns and generalize numerical laws without relying on memorization or prior knowledge.
Our experiments reveal that LLMs struggle with textual numerical data, often focusing on absolute values rather than capturing percentage changes and relations between data points, akin to algebraic reasoning. Additionally, LLMs tend to overemphasize recent data while overlooking earlier information, even when explicitly highlighted in the prompt. These limitations suggest that LLMs lack the ability to abstract numerical information into higher-level representations, a crucial skill for generalizing beyond explicit values.
In contrast, LLMs perform better when presented with numerical data in visualized formats, such as scatter plots, line charts, and bar graphs, which involve geometric reasoning. Our experiments show that LLMs processing visual numerical data consistently outperform those processing textual numerical data. As shown in Figure 1, visual numerical data enables LLMs to achieve significantly higher return rates in the Agent Trading Arena, underscoring the advantages of structured visual representations.
Incorporating the reflection module (Zhou et al., 2022; Shinn et al., 2023) further emphasizes the performance gap between textual and visual data, with the module proving particularly effective for visual representations. By leveraging structured visualizations, the reflection module enhances the model’s ability to reason through complex data relations, leading to more accurate and strategic decision-making.
The Agent Trading Arena provides a valuable framework for advancing LLM generalization and adaptability, with broad applications in fields like finance, healthcare, and scientific research. To further validate our findings, we extend our evaluation to real-world financial data, such as NASDAQ Stock dataset, which includes historical trading data. The consistent results suggest an important contribution: LLMs exhibit stronger reasoning with visual geometric data compared to textual numerical data, demonstrating the value of visual representations in enhancing LLM performance.
@@ 2 Related Works
2.1 Mathematical Benchmarks for LLMs
Math word problems (MWPs) have been widely studied, leading to the development of various benchmarks for evaluating models’ mathematical reasoning and problem-solving abilities. Early datasets, such as MAWPS (Koncel-Kedziorski et al., 2016), standardized existing problems to facilitate consistent evaluation. Math23K (Wang et al., 2017) introduced a large-scale collection of Chinese arithmetic problems that require structured equation solving. To increase diversity, benchmarks like ASDiv (Miao et al., 2020) and SVAMP (Patel et al., 2021) provide richer annotations and a broader range of problem types. More recent benchmarks, including GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021), focus on multi-step reasoning and advanced mathematical concepts, broadening the scope of evaluation. Additionally, MathQA-Python (Austin et al., 2021), a Python variant of MathQA (Amini et al., 2019), emphasizes programmatic reasoning, while MGSM (Shi et al., 2023) extends these benchmarks to multilingual contexts. Despite these advancements, current models (Lu et al., 2023b, a) primarily rely on memory-based answering strategies learned through extensive training, rather than demonstrating true mathematical reasoning.
2.2 LLMs for Enhanced Mathematical Reasoning
Building on these benchmarks, LLMs have advanced mathematical problem-solving by incorporating specialized datasets into their training. Models such as Galactica (Taylor et al., 2022), PaLM-2 (Anil et al., 2023), Minerva (Lewkowycz et al., 2022), and LLaMA-2 (Touvron et al., 2023) leverage extensive datasets during pre-training, improving their mathematical reasoning abilities and understanding of complex concepts. Fine-tuned models like MetaMath (Yu et al., 2024), MAmmoTH (Yue et al., 2024), and WizardMath (Luo et al., 2023) are specifically tailored for mathematical tasks. These models undergo domain-specific fine-tuning with carefully curated datasets, enabling them to tackle advanced reasoning challenges and significantly enhance performance on specialized benchmarks.
However, existing methods (Naveed et al., 2023; Yang et al., 2024a) often rely on large training datasets, raising concerns about the true reasoning capabilities of LLMs. While these methods primarily assess performance on established problem types, the heavy reliance on extensive training data suggests that models may achieve high performance through memorization and pattern recognition, rather than genuine reasoning. Consequently, alternative evaluation paradigms are needed to more accurately assess LLMs’ ability to generalize mathematical principles to novel scenarios.
Figure 2: Stock Trading Workflow in Agent Trading Arena. Top: Workflow of a trading day, including preparation, trading, and post-trading reflection. Agents discuss insights in the chat pool, analyze market trends, execute trades, and refine strategies based on performance. Bottom: Example of agents’ interactions in the chat pool and dynamic strategy updates.
@@ 3 Proposed Method
To mitigate the influence of human prior knowledge and memory, we designed a closed-loop economic system (Guo et al., 2024) called the Agent Trading Arena, a zero-sum game simulating complex, quantitative real-world scenarios. The simulation workflow is illustrated in Figure 2 and further detailed in Appendix A. In the Agent Trading Arena, agents can invest in assets, earn dividends from holding assets, and pay daily expenses using virtual currency. The agent with the highest total return wins the game.
3.1 Agent Trading Arena
Structure of Agent Trading Arena.
To eliminate external knowledge biases, asset prices are determined by a bid-ask system, reflecting the prices at which buyers and sellers are willing to transact. The system evolves solely based on agents’ actions and interactions, without external influences. This design ensures that the outcomes of agents’ actions are not immediately apparent but unfold gradually, influenced by other agents’ decisions.
To encourage active participation, a dividend mechanism is introduced. There are two primary sources of income in this system: capital gains from asset price differentials and dividends from holding assets. Dividends for each asset are distributed according to a predefined ratio, serving as an implicit anchor for asset prices. Agents holding more low-cost assets receive higher dividends. To prevent passive asset holding until the end of the game, agents must pay a daily capital cost proportional to their total wealth. These expenses are offset by asset dividends, and only agents with sufficient low-cost assets can cover costs. Under the pressure of significant daily expenses, agents must act swiftly and strategically, triggering frequent trades and price fluctuations to stimulate market activity. This dynamic mechanism ensures fairness in the zero-sum game while preventing agents from relying on fixed strategies to find optimal solutions.
Agents Learn and Compete in Arena.
The zero-sum game structure is crucial to eliminating the possibility of a universally optimal strategy. In fixed scenarios with a static optimal solution, agents could rely on predefined rules or memory-based approaches, bypassing adaptive decision-making. The zero-sum game ensures that there is no universally correct solution, with outcomes evolving dynamically based on agent interactions and competition. This design forces agents to continually adapt, learn from feedback, and develop context-dependent strategies, promoting deeper environmental exploration and preventing reliance on static or memory-driven solutions.
In the Agent Trading Arena, agents are unaware of implicit rules, except for the objective to maximize their virtual wealth throughout the simulation. To win this zero-sum game, agents must effectively learn from experience, decipher hidden game rules, and develop strategies to counter competitors. This requires the ability to comprehend numerical feedback, formulate enduring strategies, and make informed decisions. Unlike other mathematical reasoning problems, the results of their actions unfold gradually and dynamically. Moreover, agents are easily misled by erroneous information from competitors, hindering their ability to discern strategic cues from competitors’ textual data. Importantly, agents remain unaware of these implicit rules, so applying real-world knowledge does not benefit their performance. Therefore, agents must rely on experiential learning to decipher the hidden game rules and ultimately achieve victory.
3.2 Types of Numerical Data Input
Limitations of Textual Numerical Data.
In the Agent Trading Arena, the generated stock data is stored in numerical format. When used directly as input to an LLM, the models often struggle to interpret numerical data accurately or make sound decisions. To mitigate this, we convert the data into textual formats (Hu et al., 2024; Zhang et al., 2024), enhancing semantic features and clarifying output requirements to improve the models’ understanding. During interactions, the LLMs process stock prices, trading volumes, and market indices presented as textual numerical data.
@@ 3.2 Types of Numerical Data Input
Figure 3: Textual and Visual Representations of Corresponding Inputs and Outputs. The left images display the agent’s Buy and Sell trading records, daily trade prices, and K-line charts for three stocks. The output from visual inputs (bottom right) captures overall stock trends and long-term behavior, while the output from textual inputs (top right) focuses on specific current prices.
However, this textual approach reveals significant limitations. While the data is presented clearly, LLMs tend to focus excessively on specific values rather than identifying long-term trends or global patterns. They also struggle with understanding correlative relations and percentage changes, limiting their ability to assess differences and identify connections between data points. When analyzing time-series data with complex patterns, LLMs often fixate on individual data points, overlooking overarching relations. This issue is evident in the analysis output in the top-right corner of Figure 3, where LLMs’ focus on individual values impedes their ability to generalize, reducing their capacity to extract meaningful global insights.
Additionally, LLMs often overemphasize recent data while undervaluing historical information, even when prompted to consider its importance. This prevents them from effectively integrating past data and recognizing long-term patterns, complicating their understanding of numerical relations and trends. These challenges highlight the need for improved mechanisms to process numerical relations, identify global trends, and derive deeper insights from textual numerical data.
Potential of Visual Numerical Data.
Since textual numerical data often leads LLMs to focus on local details while neglecting broader relations, we investigated whether visual representations, such as scatter plots, line charts, and bar charts, could help LLMs better understand overall trends, similar to human reasoning. Thus, we transition from textual numerical data inputs to visualized formats  (Yang et al., 2025). As demonstrated in the bottom-right corner of Figure 3, visual representations enable LLMs to more effectively grasp global trends, patterns, and relations that are often difficult to discern from textual numerical data alone.
These findings highlight the advantages of structured, visual numerical data, indicating that this format allows LLMs to more intuitively and comprehensively understand complex data, better capturing overall fluctuations, whereas text tends to focus on local details. By combining visualization and textual representations, LLMs not only overcome the challenges of relations in time-series data but also demonstrate better performance in identifying long-term trends and global patterns, while still attending to local details.
3.3 Reflection Module
We propose a strategy distillation method, illustrated in Figure 4, that delivers real-time feedback to LLMs by analyzing both descriptive textual and visual numerical data. This enables the generation of new strategies and optimization of action plans. The approach allows agents to evaluate their results, refine strategies, and adapt continuously based on feedback. The process begins with assessing the day’s trajectory memory and associated strategies using an evaluation function. The strategic generation process leverages contrastive analysis of peak and nadir performers from the evaluation phase, creating bidirectional learning signals that inform subsequent iterations. This iterative cycle ensures continuous strategy evolution, fostering sustained improvement in decision-making.
Figure 4: Design of the Reflection Module. The process evaluates daily trajectory memory and strategies (top right), then generates new strategies (center) based on evaluation, environmental feedback (bottom right), and feedback from the 5 top- and bottom-performing strategies. Stock visualization (bottom left) enhances reflection, driving continuous improvement.
The reflection module plays a crucial role in refining strategies by offering real-time feedback. It analyzes both descriptive textual and visual numerical data to generate new strategies and optimize action plans. Within the Agent Trading Arena, the reflection module is triggered regularly to consolidate daily trading records and evaluate the effectiveness of strategies, refining both successful and unsuccessful experiences to guide future decisions. Ineffective strategies are stored in a strategy library for future reference, allowing agents to review and learn from past experiences. Further details can be found in Appendix A.
ID Name Year Cash Occupation

### v2

# Agent Trading Arena: A Study on Numerical Understanding in LLM-Based Agents

@@ Abstract
Large language models (LLMs) have demonstrated remarkable capabilities in natural language tasks, yet their performance in dynamic, real-world financial environments remains underexplored. Existing approaches are limited to historical backtesting, where trading actions cannot influence market prices and agents train only on static data. To address this limitation, we present the Agent Trading Arena, a virtual zero-sum stock market in which LLM-based agents engage in competitive multi-agent trading and directly impact price dynamics. By simulating realistic bid-ask interactions, our platform enables training in scenarios that closely mirror live markets, thereby narrowing the gap between training and evaluation. Experiments reveal that LLMs struggle with numerical reasoning when given plain-text data, often overfitting to local patterns and recent values. In contrast, chart-based visualizations significantly enhance both numerical reasoning and trading performance. Furthermore, incorporating a reflection module yields additional improvements, especially with visual inputs. Evaluations on NASDAQ and CSI datasets demonstrate the superiority of our method, particularly under high volatility. All code and data are available at https://github.com/wekjsdvnm/Agent-Trading-Arena.
Agent Trading Arena: A Study on Numerical Understanding in LLM-Based Agents
Tianmi Ma1,#\#, Jiawei Du2,#\#, Wenxin Huang3, ✉, Wenjie Wang4, Liang Xie1, Xian Zhong1, ✉, and Joey Tianyi Zhou2 1 Wuhan University of Technology  2 Agency for Science, Technology and Research, Singapore 3 Hubei University  4 University of Science and Technology of China Correspondence: wenxinhuang_wh@163.com, zhongx@whut.edu.cn
footnotetext: #\# denotes equal contribution.
@@ 1 Introduction
In recent years, large language models (LLMs) have demonstrated strong capabilities in contextual understanding and generation across tasks such as natural language processing (Koshkin et al., 2024), data analysis, and reasoning (Olausson et al., 2023; Qiao et al., 2023). Their applications are rapidly extending to complex decision-making and interactive scenarios. LLM-based autonomous agents have emerged as a promising paradigm, showing notable progress in communication (Qian et al., 2024; Zhong et al., 2024), social simulation (Park et al., 2023, 2022; Lin et al., 2023), gaming (Wang et al., 2024a; Zhu et al., 2023), and multi-agent collaboration (Chen et al., 2023; Hu et al., 2023; Hong et al., 2024). These agents are increasingly regarded as a pathway toward general artificial intelligence, with growing potential in dynamic decision-making and interactive environments.
Figure 1: Comparison between Static Backtesting and the Arena. Left: Static backtesting performs one-shot analysis on historical data with fixed trading decisions, leveraging models such as Agent (market reasoning), StockFormer (predictive coding + RL hybrid), and StockMixer (time-series prediction). Right: The Arena, by contrast, supports real-time sensing, continuous adaptation, and feedback-driven execution for dynamic trading.
Despite these advances, the application of LLM-based agents in financial markets remains limited. Financial ecosystems naturally form complex multi-agent environments in which participants interact through dynamic market mechanisms. However, most existing studies (Cheng and Chin, 2024; Xiao et al., 2024) rely on historical backtesting, as illustrated in Fig.˜1, where agents cannot influence market prices and thus fail to capture interactivity and adaptive behavior. This gap underscores the need for a framework that supports real-time agent interactions and realistic market feedback.
To address these limitations, we present the Agent Trading Arena, a virtual zero-sum stock market in which LLM-based agents engage in competitive trading and directly affect price dynamics through bid-ask interactions. Our platform generates continuously evolving numerical data, forcing agents to adapt to shifting market conditions; any temporarily optimal strategy is quickly countered, rewarding adaptability and robust reasoning over memorization.
By simulating bid-ask order matching in real time, the Arena faithfully reproduces liquidity dynamics and price impact, enabling agents to experience authentic market friction and slippage. Unlike static backtests, the Arena continuously updates its state based on agents’ collective actions, creating an ever-shifting landscape that prevents overfitting to historical patterns. Because agents directly influence subsequent price trajectories, the Arena closes the loop between training and evaluation, effectively bridging the “train-test” gap of traditional financial simulations. Its modular design further supports seamless integration of new market features, such as varying transaction costs, order types, or macroeconomic shocks, making it a versatile testbed for financial AI research.
Our experiments show that LLMs struggle with numerical reasoning when processing plain-text data: they focus on absolute values, overlook percentage changes and relational patterns, and overemphasize recent trends even when earlier data are emphasized. In contrast, visual representations, such as line charts, bar graphs, and scatter plots, substantially improve both numerical reasoning and trading performance. As shown in Table˜5, chart-based inputs lead to significantly higher return rates in the Agent Trading Arena.
We further integrate a reflection module (Zhou et al., 2022; Shinn et al., 2023), which amplifies the benefits of visual data by enabling agents to iteratively refine their reasoning. This module is particularly effective with structured visual inputs, facilitating deeper pattern recognition and more strategic decision-making.
Finally, we validate our approach on real-world datasets from NASDAQ and CSI, demonstrating consistent performance gains over baselines, particularly under high-volatility market regimes.
Our contributions are summarized threefold:
  • We design the Agent Trading Arena, a dynamic and interactive simulation framework that moves beyond static backtesting with zero-sum multi-agent competition, enabling realistic evaluation of LLMs’ reasoning and adaptability.
  • We uncover LLMs’ limitations in textual numerical reasoning, highlighting difficulties with relational patterns, percentage changes, and long-range dependencies, while demonstrating superior performance with structured visual data.
  • We demonstrate that a reflection module enhances LLMs’ reasoning over complex visual inputs, leading to more accurate and strategic trading decisions.
@@ 2 Related Works
2.1 LLM-Based Agents in Dynamic Decision-Making Environments
LLMs serve as the core reasoning engines in autonomous agents, enabling planning, tool use, reflection, and adaptation within dynamic environments. Advances in NLP have empowered models such as ChatGPT (OpenAI, 2023), PaLM (Anil et al., 2023), and LLaMA (Touvron et al., 2023) to operate via natural language, making them particularly effective in interactive, text-based scenarios. Frameworks like ChatDev (Qian et al., 2024), AgentVerse (Chen et al., 2023), and MetaGPT (Hong et al., 2024) demonstrate multi-agent collaboration on software development tasks, while Voyager (Wang et al., 2024a), GROOT (Cai et al., 2024), and Optimus-1 (Li et al., 2024b) showcase autonomous exploration in open-ended environments such as Minecraft. Social simulators, including Generative Agents (Park et al., 2023) and AgentSims (Lin et al., 2023), further reveal emergent behaviors through sustained multi-agent interactions.
Together, these studies demonstrate that LLM-based agents can effectively coordinate, communicate, and make decisions in complex, evolving environments. Their adaptability and collaborative strengths highlight strong potential for real-time decision-making in finance, an information-dense and competitive domain analogous to the dynamic settings where LLMs excel.
2.2 LLMs in Financial Analysis and Decision Making
LLMs have been applied to sentiment analysis, financial question answering, and investment recommendation. Benchmarks such as SocioDojo (Cheng and Chin, 2024) evaluate financial understanding but lack interactive components. Multi-agent systems like StockAgent (Zhang et al., 2024a), TradingAgents (Xiao et al., 2024), and CryptoTrade (Li et al., 2024a) enhance performance through role specialization and iterative reflection. In contrast, traditional forecasting models, e.g., StockFormer (Gao et al., 2023), TimesNet (Wu et al., 2023), PatchTST (Nie et al., 2023), PDFormer (Jiang et al., 2023), StockMixer (Fan and Shen, 2024), and iTransformer (Liu et al., 2024b), focus on price prediction using advanced methods such as wavelet transforms and temporal modeling to capture market dynamics.
However, neither paradigm addresses the “live market” paradox: without real-time adaptive mechanisms, these systems cannot simulate how participants adjust strategies in response to multi-agent interactions, news shocks, and evolving order flows. This limitation significantly reduces their practical utility in real trading environments.
Figure 2: Overall architecture of the Agent Trading Arena, consisting of three components: the Chat Pool (left) for inter-agent communication, agent modules (center) for analysis, decision-making, and reflection, and the trading platform (right), where agents’ buy/sell actions dynamically determine asset prices in a closed-loop market simulation.
3 Agent Trading Arena
To overcome the limitations of static backtesting and capture dynamic interactions among market participants, we propose the Agent Trading Arena, a closed-loop, zero-sum economic system that simulates complex quantitative scenarios in real markets (Guo et al., 2024). The overall architecture is shown in Fig.˜2 and detailed in Appendix˜A. Within this environment, agents invest in assets, earn dividends on their holdings, and pay daily expenses using virtual currency.
3.1 Virtual Market Environment
Closed-Loop System.
The Agent Trading Arena is a closed-loop virtual economy that isolates external knowledge and avoids reliance on historical data, enabling fair evaluation of agent capabilities. Unlike traditional backtesting grounded in past market records, this sandboxed system removes real-world anchors: all assets are abstract symbols, and their prices emerge solely from agent interactions through a bid-ask mechanism. By eliminating external context, the environment compels agents to adapt strategies based on internal dynamics and opponent behavior, fostering genuine decision-making and strategic evolution.
Bid-Ask Mechanism.
Asset prices arise exclusively from a bid-ask process: each agent’s orders directly influence the market price, with no external benchmarks. This closed-loop design rigorously captures endogenous price formation, faithfully modeling liquidity constraints, order book dynamics, market depth, and slippage. Every execution reshapes the market microstructure, producing non-trivial price impacts and path-dependent outcomes that evolve with aggregate agent behavior over time.
3.2 Chat Pool
To simulate real-world information flow, we introduce a shared Chat Pool where agents post and read “news” items, analyst commentary, and social media-style opinions. By reproducing both informative signals and noisy interference, including hallucinations (Gekhman et al., 2024) and misinformation, the Chat Pool reflects how public information can shape, and at times distort, traders’ narratives and expectations. This process increases the complexity of strategic reasoning and adaptive behavior in the market.
3.3 Market Mechanisms
We incorporate two key incentives to sustain active trading and realistic dynamics: 1) Dividends & Capital Gains: Agents earn income from both price appreciation and dividends, distributed proportionally to holdings. This implicitly anchors prices and rewards low-cost positions. 2) Daily Wealth Fee: A cost proportional to each agent’s total wealth is charged daily. Only those earning sufficient dividends can cover this fee, which discourages passive holding and promotes frequent, strategic trades.
Together, these mechanisms maintain continuous market activity, encourage adaptive decision-making, and prevent reliance on static or exploitative strategies in a zero-sum environment.
3.4 Zero-Sum Competition
By design, the Arena is zero-sum: the profit of one agent is the loss of another, meaning there is no fixed “best” strategy. Competitors counter any temporarily successful tactic, compelling agents to learn implicit rules, interpret delayed numerical feedback, and continuously adapt in real time. With results unfolding gradually and misinformation potentially introduced by rivals, agents must rely on experiential learning and develop context-aware strategies through ongoing adaptation rather than predefined rules or external knowledge.
@@ 4 Proposed Method
Our agent, ArenaTrader, integrates textual and visual modalities to form a comprehensive understanding of market conditions. It employs a structured three-stage framework, market analysis, trade execution, and strategic reflection, to leverage this multi-modal representation. This design enables informed and adaptive decision-making based on data-driven insights and iterative feedback.
@@
Figure 3: Textual vs. Visual Input–Output Representations. Left: Agent buy/sell records, daily trade prices, and corresponding candlestick charts. Top right: Visual-input outputs highlight overall trends and long-term patterns. Bottom right: Textual-input outputs emphasize specific price points.
4.1 Numerical Data Representations
Textual Input.
Time-series stock data are represented as plain text, including timestamps, prices, and volumes, to provide explicit semantics (Hu et al., 2024; Zhang et al., 2024b). However, LLMs often fixate on absolute values, overlook percentage changes, and overemphasize recent entries, which limits their ability to reason across long horizons (see lower right of Fig.˜3).
Visual Input.
To complement textual data, we introduce visualizations such as line charts, bar graphs, and scatter plots (Yang et al., 2025), which highlight global trends, relational patterns, and volatility at a glance (see upper right of Fig.˜3). These visualizations are consistent with the textual input, ensuring a cohesive representation. This structured format enables LLMs to integrate local details with broader market context more effectively.
4.2 Agent Architecture
Investment Analysis.
Inspired by human trading behavior, each agent begins by analyzing market conditions before executing trades. This process involves extracting and synthesizing multi-modal signals from the Chat Pool, price histories, and fundamental indicators. By combining these heterogeneous inputs, the agent constructs a stock ranking framework based on predefined heuristics or learned criteria, allowing it to identify promising assets for portfolio allocation.
Action Decision.
Agents execute multiple trades per day, making buy, sell, or hold decisions by integrating diverse signals, including analysis reports, real-time market data, and historical performance trends. They evaluate volatility, price momentum, and portfolio risk exposure to compute optimal trade volumes under capital constraints. Orders are executed via the bid-ask mechanism, with portfolios updated accordingly. Each decision aims to maximize long-term returns while maintaining a competitive edge in the evolving virtual market.
Figure 4: Design of the Reflection Module. Left: Agent trading records, strategy evaluations, and top- and bottom-performing strategies from the strategy library. Top right: Corresponding stock chart. Bottom right: Newly generated strategy based on current information.
Strategy Optimization via Reflection.
At the end of each trading day, the reflection module distills daily experiences into bidirectional learning signals (see Fig.˜4). It first evaluates the effectiveness of trading trajectories and strategies using a scoring function. Then, by contrasting top- and bottom-performing strategies, agents generate refined action plans and archive unsuccessful tactics in a strategy library. This iterative process drives continuous strategy evolution and more robust decision-making.
5 Experimental Results
5.1 Experimental Setup
Environment.
The Agent Trading Arena is lightweight and accessible, requiring only API credits and no GPU for deployment. Details on the selection of different LLMs are provided in Appendix˜D.
ID Name
Duration
Initial
Capital
Profession